# Whitehead’s problem and condensed mathematics

### v2

@@ Abstract.
The first author was supported by a María Zambrano Fellowship at the University of Barcelona; the second author was supported by GAČR project 23-04683S and the Academy of Sciences of the Czech Republic (RVO 67985840); the third author was supported by GAČR project 23-05148S. A portion of this work was carried out while all three authors were participating in the Thematic Program on Set Theoretic Methods in Algebra, Dynamics and Geometry at the Fields Institute in spring of 2023. We thank the Fields Institute for their support and hospitality. The results of this paper were presented in detail at the Cornell Logic Seminar in Fall 2023. We would like to thank the participants of the seminar, particularly Mark Schachner and Justin Moore, for their engagement with the material and for a number of helpful remarks.
@@ 1. Introduction
Recall that an abelian group $A$ is Whitehead if $\mathrm{Ext}^{1}(A,\mathbb{Z})=0$. Around 1950, J.H.C. Whitehead asked whether every such group is free. The converse is easily seen to be true (every free abelian group $A$ satisfies $\mathrm{Ext}^{1}(A,G)=0$ for every abelian group $G$, and hence for $G=\mathbb{Z}$), and Stein showed soon after the question was posed that every countable Whitehead group is free [25]. Progress on the general question was slow, though, and the latter soon assumed the status of a major open problem in homological algebra. Shelah’s solution to the question 20 years later was completely unexpected, for Whitehead’s problem thereby became one of the first originating outside of logic or set theory to be proven independent of the $\sf ZFC$ axioms.
@@
Second, using Theorem 1.3, particularly clauses (2) and (3) thereof, as a starting point, we begin to investigate a more general connection between condensed mathematics and set theoretic forcing. We isolate a precise sense in which, given a locally compact Hausdorff space $X$, the condensed set $\underline{X}$ is simply an organized presentation of all forcing names for points in the canonical interpretations of $X$ in all possible set forcing extensions. We then exhibit a correspondence between satisfaction of certain simple formulas concerning a topological space $X$ in two settings: first, in a Boolean-valued model $V^{\mathbb{B}}$, and second in the context of the space $C(S,X)$ of continuous functions from the Stone space $S$ of $\mathbb{B}$ to $X$. This correspondence, stated in Theorem 9.1, allows us, e.g., to immediately deduce clause (3) of Theorem 1.3 from clause (2) of the same theorem.
@@
Second, we extend Theorem 1.2 from the class of discrete abelian groups to the class of locally compact abelian groups and show that $\underline{\mathrm{Ext}}^{1}(\underline{A},\underline{\mathbb{Z}})=0$ for a locally compact abelian group $A$ if and only if $A$ is projective in the category of locally compact abelian groups.
@@
Section 9 then establishes the aforementioned equivalence between the settings of $C(S,X)$ and $V^{\mathbb{B}}$ with respect to interpretations of certain formulas. We expect the results of this section to be applicable in a wide variety of settings, and derive clause (3) of Theorem 1.3 as an immediate corollary. In Section 10, we prove Theorem 1.4(1). In Section 11, we record an algebraic lemma converting the groups $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{A},\underline{\mathbb{Z}})(S)$ into a more tractable, and suggestive, form. This lemma yields an alternative proof of clauses (1) and (3) of Theorem 1.3 (and hence of Theorem 1.2), as well as clauses (2) and (3) of Theorem 1.4. In Section 12 we prove an extension of Theorem 1.2 to the class of locally compact abelian groups. We close out our work in Section 13 with a few concluding remarks and main open questions.
@@ 2. Condensed abelian groups
Given a Hausdorff space $X$, one can define a condensed set $\underline{X}$ by letting $\underline{X}(S)=C(S,X)$, i.e., $X(S)$ is the set of continuous functions from $S$ to $X$ for all $S\in\mathsf{Prof}$; we will sometimes term this $\underline{X}$ the condensation of $X$. This operation defines a functor from the category $\mathsf{Haus}$ of Hausdorff topological spaces to $\mathsf{Cond}$. When restricted to the full subcategory of $\mathsf{Haus}$ spanned by the compactly generated Hausdorff spaces, it is a fully faithful embedding (cf. [20, Proposition 1.7]).
@@
For each uncountable cardinal $\kappa$, there is an obvious forgetful functor from $\mathsf{CondAb}_{\kappa}$ to $\mathsf{Cond}_{\kappa}$, and this functor admits a left adjoint that sends a $\kappa$-condensed set $T$ to a $\kappa$-condensed abelian group denoted $\mathbb{Z}[T]$. Concretely, as noted in [20, §2], $\mathbb{Z}[T]$ is the sheafification of the functor from $\mathsf{ED}_{\kappa}$ to $\mathsf{Ab}$ that sends each $S\in\mathsf{ED}_{\kappa}$ to the free group on $T(S)$. Moreover, $\{\mathbb{Z}[\underline{S}]\mid S\in\mathsf{ED}_{\kappa}\}$ is, for any strong limit cardinal $\kappa$, a set of compact projective generators for $\mathsf{CondAb}_{\kappa}$. These facts straightforwardly propagate upward to the category $\mathsf{CondAb}$ of all condensed abelian groups.
@@ 3. Algebraic preliminaries
Proposition 3.1 (Clausen–Scholze, [20, Proposition 4.2]).
@@
As noted in [20, proof of Corollary 4.8] (and as follows from items (ii) and (iii) of its page 13),
@@
for any condensed abelian groups $T,M$ and any extremally disconnected profinite set $S$. When $S=\ast$, our general claim follows from the facts that $\mathbb{Z}[*]=\underline{\mathbb{Z}}$ and that $\underline{\mathbb{Z}}$ is the unit object for the condensed tensor product $\otimes$. To see the assertion for discrete abelian groups, recall that their category fully and faithfully embeds into $\mathsf{CondAb}$, and observe by [13, Props. 2.18 and 2.19] that if $\cdots\to P^{1}\to P^{0}\to A$ is a projective (hence free) resolution of $A$ then $\cdots\to\underline{P}^{1}\to\underline{P}^{0}\to\underline{A}$ is a projective resolution of $\underline{A}$. Writing $\mathcal{P}$ and $\underline{\mathcal{P}}$ for $\cdots\to P^{1}\to P^{0}\to 0$ and $\cdots\to\underline{P}^{1}\to\underline{P}^{0}\to 0$, respectively, we then have
@@
It is natural at this point in the discussion to declare a condensed abelian group $T$ to be Whitehead if $\underline{\mathrm{Ext}}^{1}(T,\underline{\mathbb{Z}})=0$. Observe next that, in close analogy with the classical situation, if $F$ is a free (discrete) abelian group, then $\underline{F}$ is Whitehead in $\mathsf{CondAb}$. To see this, let $\kappa$ be a cardinal, and let $F=\bigoplus_{\kappa}\mathbb{Z}$ be the free group on $\kappa$ generators; it follows then that $\underline{F}\cong\bigoplus_{\kappa}\underline{\mathbb{Z}}$, where the direct sum on the right is computed in $\mathsf{CondAb}$, from [13, Proposition 2.19]. By [13, Remark 3.10 and Lemma 4.2], we then have
@@
is a surjection. Recall that $\underline{\mathrm{Hom}(F,\mathbb{Z})}(S)$ consists of all continuous functions $\psi:S\rightarrow\mathrm{Hom}(F,\mathbb{Z})$ (and similarly for $K$), where $\mathrm{Hom}(F,\mathbb{Z})$ is given the compact-open topology (which, since $F$ and $\mathbb{Z}$ are discrete, is simply the topology $\mathrm{Hom}(F,\mathbb{Z})$ inherits as a subspace of $\prod_{F}\mathbb{Z}$ endowed with the product topology). Moreover, the map $\underline{\mathrm{Hom}(F,\mathbb{Z})}(S)\rightarrow\underline{\mathrm{Hom}(K,\mathbb{Z})}(S)$ sends a continuous $\psi:S\rightarrow\mathrm{Hom}(F,\mathbb{Z})$ to the map $\phi:S\rightarrow\mathrm{Hom}(K,\mathbb{Z})$ defined by letting $\phi(s)=\psi(s)\restriction K$ for all $s\in S$.
Therefore, to show that $\underline{A}$ is not Whitehead, it will suffice to find a profinite set $S$ and a continuous $\varphi:S\rightarrow\mathrm{Hom}(K,\mathbb{Z})$ that does not “lift pointwise” to a map from $S$ to $\mathrm{Hom}(F,\mathbb{Z})$ (i.e., for which there is no continuous $\psi:S\rightarrow\mathrm{Hom}(F,\mathbb{Z})$ for which $\psi(s)\restriction K=\varphi(s)$ for all $s\in S$). The construction of such $S$ and $\varphi$ will be the subject of the next sections.
@@ 4. Countable groups
By Theorem 4.1, since $A$ is countable, torsion-free, and nonfree, there is a finitely generated subgroup $X_{0}$ of $A$ such that its pure closure $X$ is not finitely generated. Let $Y_{0}$ be a finite subset of $B_{F}$ such that $X_{0}\subseteq\pi[F_{0}]$ where $F_{0}$ denotes the subgroup of $F$ generated by $Y_{0}$. Put $n_{0}=0$ and enumerate $\mathrm{Hom}(F_{0},\mathbb{Z})=\{\psi_{i}\mid i<\omega\}$. Finally, let $\varphi_{0}$ be the zero homomorphism. Assume that we have already defined $Y_{i}$, $F_{i}$, $n_{i}$ and $\varphi_{i}:K_{n_{i}}\rightarrow\mathbb{Z}$, for some $i<\omega$, in such a way that $K_{n_{i}}\subseteq F_{i}$. Consider any $z_{i}\in F$ such that $x_{i}:=\pi(z_{i})\in X\setminus\pi[F_{i}]$. Since $X$ is the pure closure of $X_{0}$, there is a positive integer $j_{i}>1$ such that $j_{i}x_{i}=\pi(j_{i}z_{i})\in X_{0}\subseteq\pi[F_{0}]$. It follows that there exists $y_{i}\in F_{0}$ such that $\pi(y_{i})=\pi(j_{i}z_{i})$, whence $s_{i}:=j_{i}z_{i}-y_{i}\in K=\mathrm{Ker}(\pi)$. We can thus express $s_{i}$ as $a_{i}+\sum_{n_{i}\leq k<n_{i+1}}b_{k}t_{k}$ where $a_{i}\in K_{n_{i}}$, $n_{i}<n_{i+1}<\omega$ and $b_{k}\in\mathbb{Z}$ for each $k\in[n_{i},n_{i+1})$. Let $Y_{i+1}$ be any finite subset of $B_{F}$ such that $Y_{i}\subseteq Y_{i+1}$ and $K_{n_{i+1}}\subseteq F_{i+1}$ where $F_{i+1}$ denotes the subgroup of $F$ generated by $Y_{i+1}$.
@@
Being free, $K$ is isomorphic to the direct sum of copies of $\mathbb{Z}$ indexed by $B_{K}$. We consider the canonical embedding $\nu:K\to\mathbb{Z}^{B_{K}}$, i.e. for every $t,s\in B_{K}$, $\nu(t)(t)=1$ and $\nu(t)(s)=0$ if $s\neq t$. Then $\nu$ is pure, i.e. $\nu[K]$ is pure in $\mathbb{Z}^{B_{K}}$ (in fact, $\nu$ is even an elementary embedding). Since $K$ is not pure in $F$, it follows that there does not exist any homomorphism $\mu:F\to\mathbb{Z}^{B_{K}}$ such that $\mu\restriction K=\nu$.
For each $t\in B_{K}$, let us denote by $p_{t}:\mathbb{Z}^{B_{K}}\to\mathbb{Z}$ the $t$-th canonical projection. Since any homomorphism into direct product is uniquely determined by specifying its compositions with all the canonical projections, there has to be some $t^{*}\in B_{K}$ such that $\varphi:=p_{t^{*}}\nu$ cannot be extended to a homomorphism from $F$ into $\mathbb{Z}$. One readily checks that $t^{*}$ and $\varphi$ possess the desired properties. ∎
@@
In the formulation of Theorem 4.2, we could not have hoped to find $\varphi$ with $(\varphi\restriction B_{K})^{-1}\{1\}$ finite, let alone a $\varphi$ satisfying the stronger condition (2) from Proposition 4.4. The reason is that any given countable torsion-free group $A$, being the union of a countable chain of finitely generated free subgroups, is isomorphic to a (pure) subgroup of $\mathbb{Z}^{\omega}/\mathbb{Z}^{(\omega)}$, cf. [27, Proposition 3.1] or [18, Theorem 3.3.2]. Here, $\mathbb{Z}^{(\omega)}$ denotes the direct sum of countably many copies of $\mathbb{Z}$, i.e., the subgroup of $\mathbb{Z}^{\omega}$ consisting of elements with finite support.
@@
It follows that $F$ is countable and thus free since all countable subgroups of $\mathbb{Z}^{\omega}$ are such, cf. [6, Theorems IV.2.3 and IV.2.8]. Since $\mu$ extends the map $\nu$, we see, in particular, that $p\mu$ extends the map $p\nu$ for any $p\in\mathrm{Hom}(\mathbb{Z}^{\omega},\mathbb{Z})$. This applies specifically for any $p$ which is a $\mathbb{Z}$-linear combination of the canonical projections $p_{n}:\mathbb{Z}^{\omega}\to\mathbb{Z}$, $n<\omega$. It remains to notice that any $\varphi\in\mathrm{Hom}(\mathbb{Z}^{(\omega)},\mathbb{Z})$ which is nonzero only on finitely many elements from $B$ is of the form $p\nu$ for $p$ as above.
On the other hand, since $\mathbb{Z}$ is a slender group ([6, Corollary III.2.4]), no homomorphism $\varphi:\mathbb{Z}^{(\omega)}\to\mathbb{Z}$ which is nonzero on infinitely many elements from $B$ can be extended to a homomorphism from $\mathbb{Z}^{\omega}$ into $\mathbb{Z}$. Of course, it can happen that a particular such $\varphi$ can be extended to a homomorphism from $F$ into $\mathbb{Z}$.

### v3

@@ Abstract.
The first author was supported by María Zambrano and Marie Skłodowska Curie (project 101110452: CatT) Fellowships at the University of Barcelona; the second author was supported by GAČR project 23-04683S and the Academy of Sciences of the Czech Republic (RVO 67985840); the third author was supported by GAČR project 23-05148S. A portion of this work was carried out while all three authors were participating in the Thematic Program on Set Theoretic Methods in Algebra, Dynamics and Geometry at the Fields Institute in spring of 2023. We thank the Fields Institute for their support and hospitality. The results of this paper were presented in detail at the Cornell Logic Seminar in Fall 2023. We would like to thank the participants of the seminar, particularly Mark Schachner and Justin Moore, for their engagement with the material and for a number of helpful remarks, and to thank Peter Scholze for valuable remarks on the paper as well.
@@ 1. Introduction
Recall that an abelian group $A$ is Whitehead if $\mathrm{Ext}^{1}(A,\mathbb{Z})=0$. Around 1950, J.H.C. Whitehead asked whether every such group is free. The converse is easily seen to be true (every free abelian group $A$ satisfies $\mathrm{Ext}^{1}(A,G)=0$ for every abelian group $G$, and hence for $G=\mathbb{Z}$), and Stein showed soon after the question was posed that every countable Whitehead group is free [31]. Progress on the general question was slow, though, and the latter soon assumed the status of a major open problem in homological algebra. Shelah’s solution to the question 20 years later was completely unexpected, for Whitehead’s problem thereby became one of the first originating unequivocally outside of logic or set theory to be proven independent of the $\sf ZFC$ axioms.
@@
Second, using Theorem 1.3, particularly clauses (2) and (3) thereof, as a starting point, we begin to investigate a more general connection between condensed mathematics and set theoretic forcing. We isolate a precise sense in which, given a locally compact Hausdorff space $X$, the condensed set $\underline{X}$ is simply an organized presentation of all forcing names for points in the canonical interpretations of $X$ in all possible set forcing extensions. We then exhibit a correspondence between the satisfaction of certain simple formulas concerning a topological space $X$ in two settings: first, in a Boolean-valued model $V^{\mathbb{B}}$, and second in the context of the space $C(S,X)$ of continuous functions from the Stone space $S$ of $\mathbb{B}$ to $X$. This correspondence, stated in Theorem 9.1, allows us, e.g., to immediately deduce clause (3) of Theorem 1.3 from clause (2) of the same theorem.
@@
Second, we extend Theorem 1.2 from the class of discrete abelian groups to the class of locally compact abelian groups and show that $\underline{\mathrm{Ext}}^{1}(\underline{A},\underline{\mathbb{Z}})=0$ for a locally compact abelian group $A$ if and only if $A$ is projective in the category of locally compact abelian groups. We record as well several observations by Peter Scholze showing that Theorem 1.2 does not extend to the category of solid abelian groups.
@@
Section 9 then establishes the aforementioned equivalence between the settings of $C(S,X)$ and $V^{\mathbb{B}}$ with respect to interpretations of certain formulas. We expect the results of this section to be applicable in a wide variety of contexts, and derive clause (3) of Theorem 1.3 as an immediate corollary. In Section 10, we prove Theorem 1.4(1). In Section 11, we record an algebraic lemma converting the groups $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{A},\underline{\mathbb{Z}})(S)$ into a more tractable, and suggestive, form. This lemma yields an alternative proof of clauses (1) and (3) of Theorem 1.3 (and hence of Theorem 1.2), as well as clauses (2) and (3) of Theorem 1.4. In Section 12 we prove an extension of Theorem 1.2 to the class of locally compact abelian groups, and record impediments to its extension to the class of solid abelian groups. We close out our work in Section 13 with a few concluding remarks and one main open question.
@@ 2. Condensed abelian groups
Given a Hausdorff space $X$, one can define a condensed set $\underline{X}$ by letting $\underline{X}(S)=C(S,X)$, i.e., $X(S)$ is the set of continuous functions from $S$ to $X$ for all $S\in\mathsf{Prof}$; we will sometimes term this $\underline{X}$ the condensation of $X$. This operation defines a functor from the category $\mathsf{Haus}$ of Hausdorff topological spaces to $\mathsf{Cond}$. When restricted to the full subcategory of $\mathsf{Haus}$ spanned by the compactly generated Hausdorff spaces, it is a fully faithful embedding (cf. [26, Proposition 1.7]).
@@
For each uncountable cardinal $\kappa$, there is an obvious forgetful functor from $\mathsf{CondAb}_{\kappa}$ to $\mathsf{Cond}_{\kappa}$, and this functor admits a left adjoint that sends a $\kappa$-condensed set $T$ to a $\kappa$-condensed abelian group denoted $\mathbb{Z}[T]$. Concretely, as noted in [26, §2], $\mathbb{Z}[T]$ is the sheafification of the functor from $\mathsf{ED}_{\kappa}$ to $\mathsf{Ab}$ that sends each $S\in\mathsf{ED}_{\kappa}$ to the free group on $T(S)$. Moreover, $\{\mathbb{Z}[\underline{S}]\mid S\in\mathsf{ED}_{\kappa}\}$ is, for any strong limit cardinal $\kappa$, a set of compact projective generators for $\mathsf{CondAb}_{\kappa}$. These facts straightforwardly propagate upward to the category $\mathsf{CondAb}$ of all condensed abelian groups.
@@ 3. Algebraic preliminaries
Proposition 3.1 (Clausen–Scholze, [26, Proposition 4.2]).
@@
As noted in [26, proof of Corollary 4.8] (and as follows from items (ii) and (iii) of its page 13),
@@
for any condensed abelian groups $T,M$ and any extremally disconnected profinite set $S$. When $S=\ast$, our general claim follows from the facts that $\mathbb{Z}[*]=\underline{\mathbb{Z}}$ and that $\underline{\mathbb{Z}}$ is the unit object for the condensed tensor product $\otimes$. To see the assertion for discrete abelian groups, recall that their category fully and faithfully embeds into $\mathsf{CondAb}$, and observe by [18, Props. 2.18 and 2.19] that if $\cdots\to P^{1}\to P^{0}\to A$ is a projective (hence free) resolution of $A$ then $\cdots\to\underline{P}^{1}\to\underline{P}^{0}\to\underline{A}$ is a projective resolution of $\underline{A}$. Writing $\mathcal{P}$ and $\underline{\mathcal{P}}$ for $\cdots\to P^{1}\to P^{0}\to 0$ and $\cdots\to\underline{P}^{1}\to\underline{P}^{0}\to 0$, respectively, we then have
@@
It is natural at this point in the discussion to declare a condensed abelian group $T$ to be Whitehead if $\underline{\mathrm{Ext}}^{1}(T,\underline{\mathbb{Z}})=0$. Observe next that, in close analogy with the classical situation, if $F$ is a free (discrete) abelian group, then $\underline{F}$ is Whitehead in $\mathsf{CondAb}$. To see this, let $\kappa$ be a cardinal, and let $F=\bigoplus_{\kappa}\mathbb{Z}$ be the free group on $\kappa$ generators; it follows then that $\underline{F}\cong\bigoplus_{\kappa}\underline{\mathbb{Z}}$, where the direct sum on the right is computed in $\mathsf{CondAb}$, from [18, Proposition 2.19]. By [18, Remark 3.10 and Lemma 4.2], we then have
@@
is a surjection. Recall that $\underline{\mathrm{Hom}(F,\mathbb{Z})}(S)$ consists of all continuous functions $\psi:S\rightarrow\mathrm{Hom}(F,\mathbb{Z})$ (and similarly for $K$), where $\mathrm{Hom}(F,\mathbb{Z})$ is given the compact-open topology (which, since $F$ and $\mathbb{Z}$ are discrete, is simply the topology $\mathrm{Hom}(F,\mathbb{Z})$ inherits as a closed subspace of $\prod_{F}\mathbb{Z}$ endowed with the product topology). Moreover, the map $\underline{\mathrm{Hom}(F,\mathbb{Z})}(S)\rightarrow\underline{\mathrm{Hom}(K,\mathbb{Z})}(S)$ sends a continuous $\psi:S\rightarrow\mathrm{Hom}(F,\mathbb{Z})$ to the map $\phi:S\rightarrow\mathrm{Hom}(K,\mathbb{Z})$ defined by letting $\phi(s)=\psi(s)\restriction K$ for all $s\in S$.
Therefore, to show that $\underline{A}$ is not Whitehead, it will suffice to find a profinite set $S$ and a continuous $\varphi:S\rightarrow\mathrm{Hom}(K,\mathbb{Z})$ that does not “lift pointwise” to a map from $S$ to $\mathrm{Hom}(F,\mathbb{Z})$ (i.e., for which there is no continuous $\psi:S\rightarrow\mathrm{Hom}(F,\mathbb{Z})$ for which $\psi(s)\restriction K=\varphi(s)$ for all $s\in S$). The construction of such an $S$ and $\varphi$ will be the subject of the next sections.
@@ 4. Countable groups
By Theorem 4.1, since $A$ is countable, torsion-free, and nonfree, there is a finitely generated subgroup $X_{0}$ of $A$ such that its pure closure $X$ is not finitely generated. Let $Y_{0}$ be a finite subset of $B_{F}$ such that $X_{0}\subseteq\pi[F_{0}]$ where $F_{0}$ denotes the subgroup of $F$ generated by $Y_{0}$. Put $n_{0}=0$ and enumerate $\mathrm{Hom}(F_{0},\mathbb{Z})=\{\psi_{i}\mid i<\omega\}$. Finally, let $\varphi_{0}$ be the zero homomorphism.
Assume that we have already defined $Y_{i}$, $F_{i}$, $n_{i}$ and $\varphi_{i}:K_{n_{i}}\rightarrow\mathbb{Z}$, for some $i<\omega$, in such a way that $K_{n_{i}}\subseteq F_{i}$. Consider any $z_{i}\in F$ such that $x_{i}:=\pi(z_{i})\in X\setminus\pi[F_{i}]$.
Since $X$ is the pure closure of $X_{0}$, there is a positive integer $j_{i}>1$ such that $j_{i}x_{i}=\pi(j_{i}z_{i})\in X_{0}\subseteq\pi[F_{0}]$. It follows that there exists $y_{i}\in F_{0}$ such that $\pi(y_{i})=\pi(j_{i}z_{i})$, whence $s_{i}:=j_{i}z_{i}-y_{i}\in K=\mathrm{Ker}(\pi)$. We can thus express $s_{i}$ as $a_{i}+\sum_{n_{i}\leq k<n_{i+1}}b_{k}t_{k}$ where $a_{i}\in K_{n_{i}}$, $n_{i}<n_{i+1}<\omega$ and $b_{k}\in\mathbb{Z}$ for each $k\in[n_{i},n_{i+1})$. Let $Y_{i+1}$ be any finite subset of $B_{F}$ such that $Y_{i}\subseteq Y_{i+1}$ and $K_{n_{i+1}}\subseteq F_{i+1}$ where $F_{i+1}$ denotes the subgroup of $F$ generated by $Y_{i+1}$.
@@
Being free, $K$ is isomorphic to the direct sum of copies of $\mathbb{Z}$ indexed by $B_{K}$. We consider the canonical embedding $\nu:K\to\mathbb{Z}^{B_{K}}$, i.e., for every $t,s\in B_{K}$, $\nu(t)(t)=1$ and $\nu(t)(s)=0$ if $s\neq t$. Then $\nu$ is pure, i.e., $\nu[K]$ is pure in $\mathbb{Z}^{B_{K}}$ (in fact, $\nu$ is even an elementary embedding). Since $K$ is not pure in $F$, it follows that there does not exist any homomorphism $\mu:F\to\mathbb{Z}^{B_{K}}$ such that $\mu\restriction K=\nu$.
For each $t\in B_{K}$, let us denote by $p_{t}:\mathbb{Z}^{B_{K}}\to\mathbb{Z}$ the $t$-th canonical projection. Since any homomorphism into a direct product is uniquely determined by specifying its compositions with all the canonical projections, there has to be some $t^{*}\in B_{K}$ such that $\varphi:=p_{t^{*}}\nu$ cannot be extended to a homomorphism from $F$ into $\mathbb{Z}$. One readily checks that $t^{*}$ and $\varphi$ possess the desired properties. ∎
@@
In the formulation of Theorem 4.2, we could not have hoped to find $\varphi$ with $(\varphi\restriction B_{K})^{-1}\{1\}$ finite, let alone a $\varphi$ satisfying the stronger condition (2) from Proposition 4.4. The reason is that any given countable torsion-free group $A$, being the union of a countable chain of finitely generated free subgroups, is isomorphic to a (pure) subgroup of $\mathbb{Z}^{\omega}/\mathbb{Z}^{(\omega)}$, cf. [34, Proposition 3.1] or [24, Theorem 3.3.2]. Here, $\mathbb{Z}^{(\omega)}$ denotes the direct sum of countably many copies of $\mathbb{Z}$, i.e., the subgroup of $\mathbb{Z}^{\omega}$ consisting of elements with finite support.
@@
It follows that $F$ is countable and thus free since all countable subgroups of $\mathbb{Z}^{\omega}$ are; see [10, Theorems IV.2.3 and IV.2.8]. Since $\mu$ extends the map $\nu$, we see, in particular, that $p\mu$ extends the map $p\nu$ for any $p\in\mathrm{Hom}(\mathbb{Z}^{\omega},\mathbb{Z})$. This applies specifically for any $p$ which is a $\mathbb{Z}$-linear combination of the canonical projections $p_{n}:\mathbb{Z}^{\omega}\to\mathbb{Z}$, $n<\omega$. It remains to notice that any $\varphi\in\mathrm{Hom}(\mathbb{Z}^{(\omega)},\mathbb{Z})$ which is nonzero only on finitely many elements from $B$ is of the form $p\nu$ for $p$ as above.
On the other hand, since $\mathbb{Z}$ is a slender group ([10, Corollary III.2.4]), no homomorphism $\varphi:\mathbb{Z}^{(\omega)}\to\mathbb{Z}$ which is nonzero on infinitely many elements from $B$ can be extended to a homomorphism from $\mathbb{Z}^{\omega}$ into $\mathbb{Z}$. Of course, it can happen that a particular such $\varphi$ can be extended to a homomorphism from $F$ into $\mathbb{Z}$.